OpenStudy (anonymous):

Find a such that the solution set of ax+3=15 is {-3} Please show your work.

OpenStudy (anonymous):

a(-3) + 3 = 15 -3a = 12 a = -4

OpenStudy (anonymous):

how did you now to replace the x with (-3)??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

hmm it says -3 is the solution set so i can put -3 instead of x for f(x) = 0 x = -3 where f(x)=ax + 3 - 15